# Sparse Portfolio Selection via Topological Data Analysis based Clustering

### v1

@@ 6.2 Analyzing the assets selected for index tracking
It is essential to understand the characteristics of assets that are selected based on the similarity matrices $K_{1}-K_{7}$ for the index tracking problem. Specifically, we focus our study on the standard deviation, VaR, and CVaR of the difference return series, that is, the asset return series minus the index return series. Additionally, we analyze downside deviation from the index, and the correlation with the index. [^21] VaR of the differenced series is the VaR relative to the index, which is the largest value by which the portfolio return can miss the index target in $1-\alpha$ fraction of cases. Similarly, CVaR relative to an index shows the mean deviation relative to the benchmark in the worst $\alpha$ cases. Our analysis focuses on the top 60 stocks, identified as the most similar to the index, in each of the 119 windows for all seven similarity matrices. The heat maps in Figures (6(a)-6(e)) illustrate the average values of these measures across all windows, providing insights into the properties of the shortlisted stocks.
@@
There are several observations. First, the assets selected using the similarity matrices $K_{6}$ and $K_{5}$ have a higher correlation than TDA-based similarity matrices, with $K_{6}$ achieving the highest correlation, which is not surprising given that the distance measure in this case is based on Pearson’s correlation. On the other hand, among TDA-based similarity matrices, $K_{2}$ and $K_{4}$ filter assets have a higher correlation with the index than $K_{1}$, $K_{3}$. Second, in terms of standard deviation and downside deviation from the index, the similarity matrices $K_{2}$ and $K_{7}$ have the lowest values among all kernels. Third, the assets filtered using the similarity matrices $K_{2}$ and $K_{7}$ have the lowest values of the VaR and CVaR of the difference series (assets returns and index returns). Other kernels show varying behavior in different windows, but on average, assets picked using $K_{3}$, $K_{5}$, and $K_{6}$ have a higher tail deviation from the index. In conclusion, the assets selected using $K_{2}$ are most suitable for the index tracking problem, as further confirmed in Section 5. The robustness of the kernel $K_{2}$ is established from Figures 7(a)-7(e), which shows similar plots but for a randomly chosen window during the COVID-19 period.
@@
Figure 6: Average values of five measures (correlation, standard deviation, downside deviation, VaR, CVaR) across all windows for the top 60 stocks (rows) and for the similarity matrices $K_{1}-K_{7}$ (columns)
@@
Figure 7: Values of five measures (correlation, standard deviation, downside deviation, VaR, CVaR) for a randomly selected window during the COVID-19 period. The values are plotted for the top 60 stocks (rows) and for the similarity matrices $K_{1}-K_{7}$ (columns)

### v2

@@ Appendix C Analyzing the assets selected for index tracking
It is essential to understand the characteristics of assets that are selected based on the similarity matrices $K_{1}-K_{7}$ for the index tracking problem. Specifically, we focus our study on the standard deviation, VaR, and CVaR of the difference return series, that is, the asset return series minus the index return series. Additionally, we analyze downside deviation from the index, and the correlation with the index. [^28] VaR of the differenced series is the VaR relative to the index, which is the largest value by which the portfolio return can miss the index target in $1-\alpha$ fraction of cases. Similarly, CVaR relative to an index shows the mean deviation relative to the benchmark in the worst $\alpha$ cases. Our analysis focuses on the top 60 stocks, identified as the most similar to the index, in each of the 119 windows for all seven similarity matrices. The heat maps in Figures (3(a)-3(e)) illustrate the average values of these measures across all windows, providing insights into the properties of the shortlisted stocks.
@@
There are several observations. First, the assets selected using the similarity matrices $K_{6}$ and $K_{5}$ have a higher correlation than TDA-based similarity matrices, with $K_{6}$ achieving the highest correlation, which is not surprising given that the distance measure in this case is based on Pearson’s correlation. On the other hand, among TDA-based similarity matrices, $K_{2}$ and $K_{4}$ filter assets have a higher correlation with the index than $K_{1}$, $K_{3}$. Second, in terms of standard deviation and downside deviation from the index, the similarity matrices $K_{2}$ and $K_{7}$ have the lowest values among all kernels. Third, the assets filtered using the similarity matrices $K_{2}$ and $K_{7}$ have the lowest values of the VaR and CVaR of the difference series (assets returns and index returns). Other kernels show varying behavior in different windows, but on average, assets picked using $K_{3}$, $K_{5}$, and $K_{6}$ have a higher tail deviation from the index. In conclusion, the assets selected using $K_{2}$ are most suitable for the index tracking problem, as further confirmed in Section 5. The robustness of the kernel $K_{2}$ is established from Figures 4(a)-4(e), which shows similar plots but for a randomly chosen window during the COVID-19 period.
@@
Figure A.3: Average values of five measures (correlation, standard deviation, downside deviation, VaR, CVaR) across all windows for the top 60 stocks (rows) and for the similarity matrices $K_{1}-K_{7}$ (columns)
@@
Figure A.4: Values of five measures (correlation, standard deviation, downside deviation, VaR, CVaR) for a randomly selected window during the COVID-19 period. The values are plotted for the top 60 stocks (rows) and for the similarity matrices $K_{1}-K_{7}$ (columns)